# DISCO: Diversifying Sample Condensation for Efficient Model Evaluation

### v1

@@ Appendix C Vision results
Model Split. As in the language domain, we use the chronological split. The cutoff date is 5 April 2023. The train-test ratio of models is 88:12.
Evaluation. We use mean absolute error (MAE) and Spearman rank correlation between the true and predicted performances.
Results. Table 3 shows the main results. Uniform random sampling, together with direct evaluation with the corresponding annotated labels, yields 3.03%p MAE and .652 rank correlation at 100 samples.
@@
We evaluate the effectiveness of DISCO in two stages. First, we apply the model-signature approach using uniform random sampling. Then, we enhance it by selecting samples based on predictive diversity score (PDS). The results follow a similar trend. With uniform random sampling, model signatures combined with Random Forest achieve 0.86%p MAE and a rank correlation of .944, significantly outperforming the naive baseline. Incorporating PDS further improves performance, reaching 0.63%p MAE and a rank correlation of .969.
To illustrate how well the estimated performances align with the true values, we present a scatter plot in Figure 6. The high Pearson correlation coefficient of .969 indicates a strong agreement between the two.
@@
Figure 7: ImageNet performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 3. Main observations: Same as for language experiments DISCO hits a better efficiency-precision trade-off across all range of compression rates.

### v2

@@ Appendix I Implementation details
Model Split. As in the language domain (§ 5.2), we use the chronological split. The cutoff date is 5 April 2023. The train-test ratio of models is 88:12.
Metrics. We use mean absolute error (MAE) and Spearman’s rank correlation between the true and predicted performances.
Evaluation. Evaluation protocol follows the one in § 5.1
J.2 About baselines for vision domain
Our work is not the first to propose efficient evaluation in the vision domain. The two closest methods are Lifelong Benchmark (Prabhu et al., 2024) (NeurIPS 2024) and SSEPY (Fogliato et al., 2024) (ECCV 2024). They propose efficient evaluation methods for visual models, using a similar two-stage framework for efficient evaluation in the language domain (see § 4 / Figure 2 of our submission):
  1. 1.
    Select “important/representative” anchor points.
  2. 2.
    Estimate model performance based on model outputs on the anchor points.
In (Prabhu et al., 2024), mean correctness scores across source models are used to measure sample difficulty, and anchor points are selected by sampling every k-th datapoint after sorting them by difficulty (where $k=\frac{\#\text{all datapoints}}{\#\text{anchor points}}$). Final performance is predicted as a weighted sum of correctness scores predicted for each test datapoint. Predicted correctness scores are binary values indicating relative position (after sorting by difficulty) to the hardest anchor point the target model got right. In (Fogliato et al., 2024), confidence scores of the target model are used to measure sample difficulty. Then samples are clustered by difficulty with K-Means, and anchor points are selected as the centroids of the clusters. Final performance is predicted as a weighted sum of anchor correctness scores. Weights for the weighted sum are determined based on the corresponding cluster sizes using the Horvitz-Thompson estimator.
The main message of our paper is that for selecting anchor points for efficient evaluation, it is better to select diversity-inducing data points (DISCO) than to make a good coverage of sample difficulty (prior work). In § 5.3, we have shown that our approach beats prior approaches in the language domain.
Likewise, in the vision domain, the existing approaches (Prabhu et al., 2024; Fogliato et al., 2024) focus on a good coverage of sample difficulty rather than on maximizing per-sample information by seeking diversity-inducing data points. We test empirically in Table 3 whether the same conclusion holds in the vision domain by comparing DISCO to (Prabhu et al., 2024; Fogliato et al., 2024).
We computed the results for these baselines ourselves, as the papers do not contain results on ImageNet. For fair comparison, we use the same setup as described in § J.1.
J.3 Main results
Table 3 shows the main results. See § 5.5 for their overview.
We evaluate the effectiveness of DISCO in two stages. First, we apply the model-signature approach using uniform random sampling. Then, we enhance it by selecting samples based on predictive diversity score (PDS). The results follow a similar trend to the language domain. With uniform random sampling, model signatures combined with Random Forest achieve 0.86%p MAE and a rank correlation of .944, significantly outperforming the naive baseline. Incorporating PDS further improves performance, reaching 0.63%p MAE and a rank correlation of .969.
@@
To illustrate how well the estimated performances align with the true values, we present a scatter plot in Figure 9. The high Pearson correlation coefficient of .970 indicates a strong agreement between the two.
@@
Figure 10: ImageNet performance estimation vs. compression rates. Mean absolute error (MAE), measured in %p difference in accuracy, and the Spearman rank correlation between the true model ranking and the estimated model ranking are shown. At 100 samples, the results are identical to Table 3. Main observations: Same as for language experiments DISCO hits a better efficiency-precision trade-off across the entire range of compression rates.